OpenStudy (anonymous):

The price of an answering machine is reduced 15%. The sales tax is 8%. Is the final cost equal to the original price less than 7%? Explain.

OpenStudy (anonymous):

The price of an answering machine is reduced 15%. The sales tax is 8%. This is saying:\[ C = P_{old}\times (1-0.15)\times(1+0.08) \] Is the final cost equal to the original price less than 7%? This is asking: \[ C = P_{old}\times (1-0.07) \]